OpenStudy (anonymous):

evaluate 12!/12! a. 0 b. 1 c. 12

OpenStudy (anonymous):

B)

jabez177 (jabez177):

@UnkownBeast911, please do not give out directs answers... What does 1/1 equal @HeatherMae2373?

OpenStudy (anonymous):

Agreed ^^^

jabez177 (jabez177):

\[\frac{ 12 }{ 12 } = \frac{ ? }{ ? }\]

OpenStudy (anonymous):

It's basic simplification.

jabez177 (jabez177):

Yes. But as OpenStudiers, we're supposed to "Guide"people to the answers...

jabez177 (jabez177):

If they do not know.

OpenStudy (anonymous):

does 12/12 =1?

jabez177 (jabez177):

 If you combine all these little sections. How many rectangles do you get?